OpenStudy (anonymous):

use analytic methods to find the extreme values of f(x)= (1/x) + lnx on the interval 0.5 ≤ x ≤ 4 and where they occur

OpenStudy (anonymous):

and i know that the derivative is f'(x) = -1/x^2 + 1/x but i dont know where to go from there... lol

OpenStudy (baldymcgee6):

Where do extreme values occur?

OpenStudy (anonymous):

i dont know thats what i need help finding.. lol. i just dont know how to find them.

OpenStudy (baldymcgee6):

okay, so in theory, extreme values will occur where the derivative of the function is equal to zero (i.e. a horizontal slope where there is maxima or minima), and they also occur where the derivative is undefined.

OpenStudy (anonymous):

okay sooo i set the derivative equal to zero...

OpenStudy (baldymcgee6):

and...?

OpenStudy (anonymous):

well i mean plugging in 1 for x would give you zero.

OpenStudy (baldymcgee6):

okay, so 1 is one of our critical points

OpenStudy (anonymous):

now what do we do?

OpenStudy (baldymcgee6):

Where is f'(x) = -1/x^2 + 1/x undefined?

OpenStudy (anonymous):

when x = 0?

OpenStudy (baldymcgee6):

exactly, but if you would notice in the question it gave us restrictions of 0.5 ≤ x ≤ 4, so we dont take 0 into account... So our critical point is 1

OpenStudy (baldymcgee6):

http://screencast.com/t/LJnPiplk

OpenStudy (anonymous):

oooh ok. so for the answer it says: max value is 1/4 + ln4 at x = 4 min value is 1 at x = 1 local max at (1/2, 2 - ln2) how did they get the max value and local max?

OpenStudy (baldymcgee6):

Well for the max value, they just put the biggest number they could, 4, into the function, they chose 4 because of the restrictions 0.5 ≤ x ≤ 4... 4 is the biggest number, i.e. giving the biggest value.

OpenStudy (baldymcgee6):

there is no local max... so i'm not sure where they got that.. might want to ask your teacher.

OpenStudy (anonymous):

oh.. ok. and how do they know that 1 is the min value? this is confusing for me ):

OpenStudy (baldymcgee6):

We already got that the minimum is at x = 1

OpenStudy (anonymous):

yea but how do you know its the minimum? :\

OpenStudy (baldymcgee6):

We know it is the minimum because at that point, the derivative = 0, this is the lowest point on the curve.

OpenStudy (anonymous):

thank you